OpenStudy (anonymous):

What is the volume of the prism? A triangular prism is shown. The triangular base has a base of seven yards and a height of five yards. The height of the prism is ten yards.

OpenStudy (anonymous):

Pleasee (:

OpenStudy (anonymous):

V=B*H B=(1/2)b*h=(1/2)(7)(5)=17.5 V=(17.5)(10)= 175

OpenStudy (anonymous):

Thank youuuu

OpenStudy (anonymous):

your welcome